# A Volumetric Approach to Privacy of Dynamical Systems

### v3

@@ 2.3 Inference Attack
We assume that the adversary observes a set of public states, denoted as $\mathcal{M}_{k|k}^{x}$, which includes the actual public state value $X_{k}=x_{k}$. Also, every element of $\mathcal{M}_{k|k}^{x}$ is considered a potential candidate for the actual public state. Notably, if the public state is directly transmitted to the adversary without any processing, the public state set $\mathcal{M}_{k|k}^{x}$ contains only the actual public state.
Based on the observed public state set, the adversary identifies all potential values of the private state that align with $\mathcal{M}_{k|k}^{x}$ to construct its uncertainty set. This process is referred to as the inference attack. We next define the inference attack recursively.
At time $k$, given the public state set $\mathcal{X}_{k-1|k-1}$ and the uncertainty private state set $\mathcal{Y}_{k-1|k-1}$, the set of public state can be predicted based on the system model (3), i.e.,
$\mathcal{X}_{k|k-1}=A_{1}\mathcal{X}_{k-1|k-1}\oplus A_{2}\mathcal{Y}_{k-1|k-1}\oplus B_{1}\mathcal{W}_{k}^{x}.$ (8)
Similarly, the set of private state can be predicted via
@@
$\mathcal{M}_{k-1|k}^{x}=$$\left\{M^{x}_{k-1|k}|X_{k|k}=A_{1}M^{x}_{k-1|k}+A_{2}Y_{k-1|k-1}+B_{1}W_{k}^{x},\right.$
$\forall X_{k|k}$ $\in\mathcal{M}^{x}_{k|k},\left.\forall Y_{k-1|k-1}\in\mathcal{Y}_{k-1|k-1},\forall W_{k}^{x}\in\mathcal{W}_{k}^{x}\right\}\!,$ (10)
$\mathcal{M}_{k-1|k}^{y}=$$\left\{M^{y}_{k-1|k}|X_{k|k}=A_{1}X_{k-1|k}+A_{2}M^{y}_{k-1|k}+B_{1}W_{k}^{x},\right.$
$\forall X_{k|k}$ $\in\mathcal{M}^{x}_{k|k},\left.\forall X_{k-1|k-1}\in\mathcal{X}_{k-1|k-1},\forall W_{k}^{x}\in\mathcal{W}_{k}^{x}\right\}\!,$ (11)
@@
where it first computes the possible sets of the public and private states, i.e., $\mathcal{M}_{k-1|k}^{x}$ and $\mathcal{M}_{k-1|k}^{y}$, based on the system model (3) and the observation $\mathcal{M}^{x}_{k|k}$ in (2.3) and (2.3), and then reduces the uncertainty sets of $X_{k-1}$ and $Y_{k-1}$ via intersection operations in (12) and (13).
@@
Finally, the public state set can be further calibrated via intersection,
@@
is the predicted uncertainty set of $X_{k}$ based on the calibrated sets $\mathcal{X}_{k-1|k}$ and $\mathcal{Y}_{k-1|k}$.
Starting from $k=0$, with the initial uncertainty sets $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, the adversary can recursively update the uncertainty sets of $X_{k}$ and $Y_{k}$ via the backward calibration (2.3)-(13), and the forward inference (14)-(16). The backward calibration (2.3)-(13) reduces the uncertainty of $X_{k-1}$ and $Y_{k-1}$, which leads to the following proposition.
@@
For any $k\geqslant 1$, the adversary’s uncertainty private state set (14) is a subset of its prediction set (9). Also, given the uncertainty sets $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$ containing the actual system states $X_{k-1}=x_{k-1}$ and $Y_{k-1}=y_{k-1}$, if the observation set $\mathcal{M}^{x}_{k|k}$ contains the actual public state $X_{k}=x_{k}$, then the inference result $\mathcal{Y}_{k|k}$ also contains the actual private state $Y_{k}=y_{k}$.
@@
The proof is omitted due to the simplicity.
According to Proposition 1, the inference attack results in a smaller uncertainty private state set $\mathcal{Y}_{k|k}$ that contains the actual private state $y_{k}$, meaning that there are fewer possible private states corresponds to the same observation set $\mathcal{M}_{k}^{x}$. In particular, if the uncertainty set $\mathcal{Y}_{k|k}$ contains only one element, then the adversary can obtain actual private state. In conclusion, the adversary can reduce its uncertainty of the private state via the inference attack.
2.4 Privacy Measure and Defense
The aforementioned inference attack relies on set operations, and its outcome is the uncertainty set of the private state, denoted as $\mathcal{Y}_{k|k}$. The first key question addressed in this paper is how to evaluate the privacy level of the private state using the uncertainty set $\mathcal{Y}_{k|k}$. Also, set operations, such as the addition in (2.3) and the intersection in (12), involve all possible combinations of elements from different collections. These operations are computationally challenging due to the continuous nature of the state space. Consequently, it is essential to provide efficient computational tools for privacy level evaluation. Moreover, from a defense perspective, another critical task is the design of effective mechanisms to mitigate privacy leakage.
To address these challenges, we propose a volumetric privacy measure and its computation approach in Section 3. Furthermore, we present an optimal privacy filter design to reduce privacy leakage in Section 4.
3 Volumetric Privacy Measure
In this section, we define the volume of uncertainty sets as a quantitative measure of the privacy level. We then introduce an interval-based approach to approximate the computation of state uncertainty sets and evaluate the corresponding privacy levels. Additionally, we analyze the properties of the volumetric measure, showing that both the privacy level and the uncertainty reduction are bounded by the new information derived from the observation set.
@@ 3.1 Privacy and Utility Measures
As discussed in Section 2.3, the uncertainty set of private state $\mathcal{Y}_{k|k}$ encompasses all possible elements that correspond to the same observation set, $\mathcal{M}_{k|k}^{x}$. However, due to the continuous nature of the state space, the number of elements in such sets is uncountable. To address this, we propose using the volume of the uncertainty set as a quantitative measure of privacy, defined as follows,
$P_{k}\left(\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right),$ (17)
where $\mathrm{Vol}\left(\cdot\right)\geqslant 0$ is a Lebesgue measure. When $\mathcal{Y}_{k|k}$ is an interval, we can compute its volume via $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)=\sum_{i=1}^{n}{\left(\overline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k}\left(i\right)\right)}$, where $\overline{Y}_{k|k}$ and $\underline{Y}_{k|k}$ are the upper and lower endpoints of the interval $\mathcal{Y}_{k|k}$, respectively.
@@
Since the uncertainty set $\mathcal{Y}_{k|k}$ contains the actual value of private state $Y_{k}=y_{k}$, specially, the adversary can accurately access to $y_{k}$ if the volume of $\mathcal{Y}_{k|k}$ is zero. On the other hand, if the volume $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ is large, then the range of private state will be large, which increases the difficulties of selecting the correct value as the estimation result. Therefore, the volume of $\mathcal{Y}_{k|k}$ describes the amount of the adversary’s inference uncertainty about the private state as well as the privacy level. We can increase the value of $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ to increase the privacy level to protect the system from inference attack.
Besides, since the uncertainty set $\mathcal{Y}_{k|k}$ is a subset of its predicted version $\mathcal{Y}_{k|k-1}$, we have
@@
$\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)\leqslant\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right),$ (18)
meaning that the adversary’s inference uncertainty about the private state is bounded by its prior knowledge $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$. In other words, the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ cannot exceed $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$.
Furthermore, we propose to use the following measure to measure the utility of public state set $\mathcal{X}_{k|k}$,
$U_{k}\left(\mathcal{X}_{k|k}\right)=\frac{1}{\mathrm{Vol}\left(\mathcal{X}_{k|k}\right)},$ (19)
since a larger $\mathcal{X}_{k|k}$ provides more possible values of the public state to the receiver, i.e., increases the estimation error. Consequently, a larger value of $U_{k}\left(\mathcal{X}_{k|k}\right)$ indicates a higher data utility of the public state set.
3.2 Inference Attack Approximation
In this subsection, we investigate the inference approximation for multi-dimensional systems using the tightest interval method [1]. This approach simplifies the computation of the privacy measure and provides the foundation for designing the optimal privacy filter as discussed in Section 4.

### v4

@@ 2.3 Inference Attack
Due to the presence of unknown disturbance terms belonging to bounded sets, multiple private states may correspond to the same public state. Consequently, for a given public state set, an adversary cannot determine the exact private state; instead, it can only identify a corresponding uncertainty set. we next define how the adversary infers the private state using set-theoretic operations.
We assume that the adversary observes a set of public states, denoted as $\mathcal{M}_{k|k}^{x}$, which includes the actual public state value $X_{k}=x_{k}$ and other elements to obfuscate the adversary’s estimation of private states. Based on the observed public state set, the adversary identifies all potential values of the private state that align with $\mathcal{M}_{k|k}^{x}$ to construct its uncertainty set. This process is referred to as the inference attack. We next define the inference attack recursively.
At time $k$, given the public state set $\mathcal{X}_{k-1|k-1}$ and the uncertainty private state set $\mathcal{Y}_{k-1|k-1}$, the set of states can be predicted based on the system model (3), i.e.,
$\mathcal{X}_{k|k-1}$ $=A_{1}\mathcal{X}_{k-1|k-1}\oplus A_{2}\mathcal{Y}_{k-1|k-1}\oplus B_{1}\mathcal{W}_{k}^{x},$ (4)
@@
$\!\!\!\!\mathcal{M}_{k\!-\!1|k}^{x}\!\!=\!A_{1}^{\!-\!1}\!\mathcal{M}_{k|k}^{x}$ $\!\oplus\!\left(\!-\!A_{1}^{\!-\!1}\!A_{2}\!\right)\!\mathcal{Y}_{k\!-\!1|k\!-\!1}\!\oplus\!\left(-\!A_{1}^{\!-\!1}\!B_{1}\!\right)\!\mathcal{W}_{k}^{x}\!,\!\!$ (6)
$\!\!\!\!\mathcal{M}_{k\!-\!1|k}^{y}\!=\!A_{2}^{\!-\!1}\!\mathcal{M}_{k|k}^{x}$ $\!\oplus\!\left(\!-A_{2}^{\!-\!1}\!A_{1}\!\right)\!\mathcal{X}_{k\!-\!1|k\!-\!1}\!\oplus\!\left(\!-A_{2}^{\!-\!1}\!B_{1}\!\right)\!\mathcal{W}_{k}^{x}\!,\!\!$ (7)
@@
where it first computes the possible sets of the public and private states, i.e., $\mathcal{M}_{k-1|k}^{x}$ and $\mathcal{M}_{k-1|k}^{y}$, based on the system model (3) and the observation $\mathcal{M}^{x}_{k|k}$ in (6) and (7), and then reduces the uncertainty sets of $X_{k-1}$ and $Y_{k-1}$ via intersection operations in (8) and (9).
@@
Finally, the public state set can be further calibrated via,
@@
where $\mathcal{M}^{x}_{k|k-1}$ is the predicted uncertainty set of $X_{k}$ based on the calibrated sets $\mathcal{X}_{k-1|k}$ and $\mathcal{Y}_{k-1|k}$.
Starting from $k=0$, with the initial uncertainty sets $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, the adversary can recursively update the uncertainty sets of $X_{k}$ and $Y_{k}$ via the backward calibration (6)-(9), and the forward inference (10)-(12). The backward calibration (6)-(9) reduces the uncertainty of $X_{k-1}$ and $Y_{k-1}$, which leads to the following proposition.
@@
For any $k\geqslant 1$, the adversary’s uncertainty set for the private state (10) is a subset of its corresponding prediction set (5), i.e., $\mathcal{Y}_{k|k}\subseteq\mathcal{Y}_{k|k-1}$. Moreover, given uncertainty sets $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$ that contain the true system states $X_{k-1}\!=\!x_{k-1}$ and $Y_{k-1}\!=\!y_{k-1}$, if the observation set $\mathcal{M}^{x}_{k|k}$ contains the true public state $X_{k}\!=\!x_{k}$, then the inference set $\mathcal{Y}_{k|k}$ contains the true private state $Y_{k}\!=\!y_{k}$.
@@
Since $\mathcal{M}^{x}_{k|k}$ contains $x_{k}$, it follows from (6) that $x_{k-1}\!\in\!\mathcal{M}^{x}_{k-1|k}$. As $x_{k-1}$ also belongs to $\mathcal{X}_{k-1|k-1}$, their intersection $\mathcal{X}_{k-1|k}$ necessarily contains $x_{k-1}$. By the same reasoning, $y_{k-1}\!\in\!\mathcal{Y}_{k-1|k}$. Propagating through the system dynamics $\mathbf{G}_{1}$ yields $y_{k}\!\in\!\mathcal{Y}_{k|k}$. Finally, since $\mathcal{X}_{k-1|k}\!\subseteq\!\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k}\!\subseteq\!\mathcal{Y}_{k-1|k-1}$, we have $\mathcal{Y}_{k|k}\!\subseteq\!\mathcal{Y}_{k|k-1}$.
According to Proposition 1, the adversary can reduce its uncertainty of the private state via the inference attack since it can obtain a smaller uncertainty private state set $\mathcal{Y}_{k|k}$ that contains the actual private state $y_{k}$. In particular, if the uncertainty set $\mathcal{Y}_{k|k}$ contains only one element, then the adversary can obtain the actual private state.
@@ 2.4 Privacy and Utility Measures
As discussed in Section 2.3, the uncertainty set of private state $\mathcal{Y}_{k|k}$ encompasses all possible elements that correspond to the same observation set, $\mathcal{M}_{k|k}^{x}$. The attacker has more inference uncertainty about the private state if $\mathcal{Y}_{k|k}$ contains more elements. However, since the state space is continuous, the number of elements in such sets is uncountable. To address this, we propose to use the volume of the uncertainty set as a quantitative measure of privacy. Specifically, we define
$\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)=\int_{\mathcal{Y}_{k|k}}\mathrm{d}y,$ (13)
where $\mathrm{Vol}(\mathcal{Y}_{k|k})$ denotes the Lebesgue measure (i.e., the geometric volume) of the set $\mathcal{Y}_{k|k}\subseteq\mathcal{R}^{n}$. We then define the privacy measure at time $k$ as
$P_{k}\left(\mathcal{Y}_{k|k}\right):=\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right).$ (14)
@@
Since the uncertainty set $\mathcal{Y}_{k|k}$ contains the actual value of private state $Y_{k}=y_{k}$, the adversary can accurately access to $y_{k}$ if the volume of $\mathcal{Y}_{k|k}$ is zero. To protect the system from inference attack, we can increase the privacy level of the private state via maximizing $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$.
On the other hand, it causes more utility distortion of the public state if the observation set $\mathcal{M}^{x}_{k|k}$ contains more elements, since it is more difficult for the receiver to recover the actual public state. To address this, we would reduce distortion via maximizing the following defined utility of public state set $\mathcal{M}^{x}_{k|k}$,
$U_{k}\left(\mathcal{M}^{x}_{k|k}\right)=1/\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right).$ (15)
Notably, set operations in the proposed inference attack and the volume computation are generally computational costly due to the continuous nature of the state space. Consequently, it is essential to provide efficient computational tools for privacy level evaluation. Also, from a defense perspective, another critical task is the design of effective mechanisms to mitigate privacy leakage. To address these challenges, we study computation approaches for inference attack in Sec. 3, and present an optimal privacy filter design to reduce privacy leakage in Sec. 4.
3 Inference Attack Approximation
As addressed in existing set-membership estimation approaches [18, 19], the set operations involved in inference attacks can be computationally expensive. To mitigate this complexity, uncertainty sets are often restricted to specific geometric forms, enabling more efficient implementation of set operations. However, more complex representations generally incur higher computational costs in volume evaluation. In the following, we present two approximation methods for implementing inference attacks and analyze their computational complexity, based on which we establish properties of volumetric privacy.
3.1 Inference Attack Approximation via CCGs
In this subsection, we show that the inference attack can be approximated using the constrained convex generator (CCG), a general set representation proposed in [30].
Definition 2.
[30] The constrained convex generator $\mathcal{Z}=\left(G,c,A,b,\mathcal{C}\right)\subset\mathcal{R}^{n}$ is defined as
$\mathcal{Z}=\left\{G\xi+c:A\xi=b,\xi\in\mathcal{C}\right\},$ (16)
where $G\in\mathcal{R}^{n\times n_{g}}$, $c\in\mathcal{R}^{n}$, $A\in\mathcal{R}^{n_{c}\times n_{g}}$, $b\in\mathcal{R}^{n_{c}}$ and $\mathcal{C}=\left\{\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{n_{p}}\right\}$, and $\mathcal{C}_{i}\subset\mathcal{R}^{m_{i}}$ are convex sets with $\sum_{i=1}^{n_{p}}{m_{i}}=n_{g}$.
The CCG encompasses a wide range of useful set representations, including zonotopes, ellipsoids, and intervals [30]. Moreover, common set operations such as the Minkowski sum and intersection admit analytical expressions, enabling its application to approximate the inference attack.
Proposition 3.
[30] Given CCGs $\mathcal{X}\!=\!\left(G_{x},c_{x},A_{x},b_{x},\mathcal{C}_{x}\right)\!\subset\!\mathcal{R}^{n}$ and $\mathcal{Y}\!=\!\left(G_{y},c_{y},A_{y},b_{y},\mathcal{C}_{y}\right)\!\subset\!\mathcal{R}^{n}$, and a matrix $R\in\mathcal{R}^{m\times n}$, we have
$R\mathcal{X}=\left(RG_{x},Rc_{x},A_{x},b_{x},\mathcal{C}_{x}\right),$
@@ 3.1 Inference Attack Approximation via CCGs
$\mathcal{X}\!\oplus\!\mathcal{Y}$ $\!=\!\left(\!\left[\begin{matrix}G_{x}&\!G_{y}\\ \end{matrix}\right]\!,c_{x}\!+\!c_{y},\mathrm{diag}\left(\!\left[\begin{matrix}A_{x}&\!A_{y}\\ \end{matrix}\right]\!\right)\!,\left[\!\begin{array}[]{c}b_{x}\\ b_{y}\\ \end{array}\!\right]\!,\left\{\mathcal{C}_{x},\mathcal{C}_{y}\right\}\!\!\right)\!,$ (19)
$\mathcal{X}\cap\mathcal{Y}$ $=\left(\left[\begin{matrix}G_{x}&0\\ \end{matrix}\right],c_{x},\left[\begin{matrix}A_{x}&0\\ 0&A_{y}\\ G_{x}&-G_{y}\\ \end{matrix}\right],\left[\begin{array}[]{c}b_{x}\\ b_{y}\\ c_{y}-c_{x}\\ \end{array}\right],\left\{\mathcal{C}_{x},\mathcal{C}_{y}\right\}\right).$ (23)
According to the computation rules in Proposition 3, the inference attack from (6) to (12) can be directly implemented, if we assume that the uncertainty sets $\mathcal{W}^{x}_{k}$, $\mathcal{W}^{y}_{k}$, $\mathcal{X}_{0|-1}$, and $\mathcal{Y}_{0|-1}$ are represented as CCGs. However, as shown in the next lemma, the computational complexity of CCG-based inference grows exponentially over time.
Proposition 4.
The computational complexity of the CCG-based inference attack at time $k$ is at least $\mathcal{O}\!\left(c^{k-1}n^{3}\right)$ for some constant $c>1$, and both the column dimension of the generator matrices $G$ and the number of constraints grow exponentially with $k$.
Proof 3.1.
The dominant operation in the inference attack from (6) to (12) is the multiplication of an $n\times n$ matrix with an $n\times m$ matrix, which has computational complexity $\mathcal{O}(mn^{2})$. For simplicity, we assume that at time $k$ the sets $\mathcal{X}_{k-1|k-1}$, $\mathcal{Y}_{k-1|k-1}$, $\mathcal{W}^{x}_{k}$, $\mathcal{W}^{y}_{k}$, and $\mathcal{M}^{x}_{k|k}$ all have generator matrices $G$ of size $n\times n$ and are described by $n$ constraints.
According to Proposition 3, after one inference step, the Minkowski sum and intersection operations cause the generator matrix in $\mathcal{X}_{k|k}$ to grow to size $n\times(c\,n)$ for some constant $c>1$, while the number of constraints increases by a factor $d>1$. Thus, both the column dimension of $G$ and the number of constraints grow exponentially with $k$. Consequently, due to the exponentially increasing column dimension, the computational complexity of matrix multiplications in the inference attack at time $k$ is at least $\mathcal{O}(c^{k-1}n^{3})$.
The high computational complexity of CCG-based inference renders real-time implementation of the inference attack and the privacy filter design in Sec. 4 intractable for large $k$. Order-reduction techniques can be employed to reduce this complexity, albeit at the cost of some loss in inference accuracy. However, such techniques also complicate the analysis of the proposed volumetric privacy metric. For clarity and focus, we defer a detailed discussion of these techniques to future work.
3.2 Inference Attack Approximation via Interval Analysis
We next consider an interval-based approximation approach for computing the inference sets. This approach is a special case of CCG-based inference but significantly reduces both computational and analytical complexity due to the efficiency of interval arithmetic. Moreover, interval-based inference serves as the foundation for designing the optimal privacy filter, as discussed in Sec. 4.
Definition 5.
The interval $\mathcal{X}=\{X\mid\underline{X}\leqslant X\leqslant\overline{X}\}$ is defined as $\begin{bmatrix}\underline{X}\\[3.00003pt] \overline{X}\end{bmatrix}$ with the lower and upper bounds $\underline{X}$ and $\overline{X}$. An interval $\mathcal{X}$ can also be expressed as a special case of the CCG representation $\mathcal{X}=\left\{\mathrm{diag}(p^{x})\,\xi+c^{x}\;:\;\xi\in\mathcal{R}^{n_{x}},\;\|\xi\|_{\infty}\leqslant 1\right\}$, with the center point $c^{x}=\frac{\overline{X}+\underline{X}}{2}$ and the radius $p^{x}=\frac{\overline{X}-\underline{X}}{2}$. Also, the volume of the interval $\mathcal{X}$ is computed as $\mathrm{Vol}\left(\mathcal{X}\right)=\sum_{i=1}^{n}{\left(\overline{X}\left(i\right)-\underline{X}\left(i\right)\right)}$ where $\overline{X}(i)$ and $\underline{X}(i)$ denote the upper and lower bounds of the $i$-th dimension of the interval $\mathcal{X}$, respectively.
Given a block matrix $A=[A_{1},\,A_{2}]$, the multiplication of $A$ with an interval $\mathcal{X}$ is defined as $A\mathcal{X}=A_{1}\underline{X}+A_{2}\overline{X}$. If $\mathcal{X}$ and $\mathcal{Y}$ are intervals, their Minkowski sum is $\mathcal{X}\oplus\mathcal{Y}=\begin{bmatrix}\underline{X}+\underline{Y}\\[3.00003pt] \overline{X}+\overline{Y}\end{bmatrix}$, and their difference, used only for volume evaluation, is $\mathcal{X}\setminus\mathcal{Y}=\begin{bmatrix}\underline{X}-\underline{Y}\\[3.00003pt] \overline{X}-\overline{Y}\end{bmatrix}$. The intersection of $\mathcal{X}$ and $\mathcal{Y}$ is $\mathcal{X}\cap\mathcal{Y}=\begin{bmatrix}\max\{\underline{X},\underline{Y}\}\\[3.00003pt] \min\{\overline{X},\overline{Y}\}\end{bmatrix}$.
We now assume that the uncertainty sets $\mathcal{W}^{x}_{k}$, $\mathcal{W}^{y}_{k}$, $\mathcal{X}_{0|-1}$, and $\mathcal{Y}_{0|-1}$ are represented as intervals. Under this assumption, the interval-based inference attack can be implemented using the following lemma.